# MindGYM: What Matters in Question Synthesis for Thinking-Centric Fine-Tuning?

### v2

@@ 5 Conclusion & Future Works
Our appendix is organized as follows:
  • Sec. A: The details of experimental setup, including the benchmarks (Sec. A.1) and the training & evaluation details (Sec. A.2).
  • Sec. B: The implementation details of MindGYM, including the description of meta-topics (Sec. B.1) and all prompts utilized in this paper (Sec. B.2).
  • Sec. C: Our exploration on multimodal data synthesis with MindGYM, including our proposed method on multimodal data (Sec. C.1), the experimental setup (Sec. C.2), experimental results (Sec. C.3) and the comprehensive analysis of the synthesis data (Sec. C.4).
  • Sec. D: The ablation stuides, including the full experimental results.
Appendix A Detailed Experimental Setup
A.1 Benchmarks
@@
Given the extensive evaluation tasks, employing the complete evaluation sets would incur substantial time expenditure. To expedite the evaluation process while ensuring fairness and accuracy, we adopt the $eval\_limit$ parameter from ms-swift, configuring GPQA to 300 samples, while retaining the original sample sizes for the remaining benchmark datasets. All experiments were conducted using this consistent setup to ensure the fairness of the experiments.
A.2 Training and Evaluation Details
@@
Appendix B Implementation Details of MindGYM
B.1 Detailed Description of Meta-topics
  • Historical Events: Analyzing past events and their consequences.
  • Economic Principles: Studying resource allocation and valuation.
  • Psychological Insights: Exploring human behavior and cognition.
  • Temporal: Reasoning across sequential or causal timelines.
B.2 Utilized Prompts
B.2.1 Stream of Consciousness
B.2.2 Prompts for Background Generation
[…]
The document should be rich enough to support multi-hop reasoning in Chinese.
B.2.3 Prompts for Seed Single-Hop Question Synthesis
[…]
Based on the background document provided, generate up to 5 logically connected sub-questions. The relationship between these sub-questions should belong to a single category:
- Bridging: requiring connecting facts or pieces of information from the document.
- Comparison: involving comparing two or more elements described in the document.
- Temporal: requiring reasoning about the order or timing of events.
Clearly state the relationship category that links all sub-questions in Chinese:
[…]
Based on the provided image, original question, and original answer, generate up to 5 logically connected sub-questions. The relationship between these sub-questions should belong to one of the following categories:
- Visual-Textual Alignment: Requiring alignment between visual (e.g., images, charts) and textual information.
- Spatial Reasoning: Involving spatial relationships or geometric layouts.
- Causal Inference: Requiring reasoning about cause-and-effect relationships.
- Contextual Synthesis: Requiring synthesis of information across multiple modalities (e.g., text, images, charts).
Clearly state the relationship category that links all sub-questions in Chinese:
B.2.4 Prompts for Challenging Multi-Hop QA Synthesis
[…]
Synthesize a multi-hop question and its answer based on the above sub-questions and background document in Chinese. Please start Qwen thinking and return the thinking process:
[…]
The question should require reasoning across the sub-questions and synthesizing information from the image and original question.
Then, provide a detailed answer to the multi-hop question, ensuring it is consistent with the image and sub-questions.
Synthesize a multi-hop question and its answer based on the above sub-questions, original image, and original question in Chinese. Please start Qwen thinking and return the thinking process:
B.2.5 Prompts for GPT-based Scorer
[…]
As a mathematical problem solver, please strictly compare and analyze the predictions of the two models:
[Problem description]
{question}
[Original model prediction]
{pred_raw}
[Fine-tuned model prediction]
{pred_finetuned}
Evaluation requirements:
Give the scores of raw and finetuned respectively
1. Reasoning depth score (0-3 points):
- 3 points: multi-step derivation with verification
- 2 points: complete derivation steps
- 1 point: simple calculation steps
- 0 points: no derivation process
2. Reasoning breadth score (0-3):
- 3 points: Explores multiple valid methods/angles, justifies optimal choice
- 2 points: Mentions alternative approaches briefly but focuses on one
- 1 points: Suggests another method without analysis
- 0 points: Single approach with no alternatives considered
3. Comparative analysis: use bullet points to list the main improvements/degradations
Please return a JSON that strictly follows this format: "accuracy_raw": 0-2, "accuracy_finetuned": 0-2, "depth_raw": 0-3, "depth_finetuned": 0-3, "comparison": ["Point 1", "Point 2"]
@@ C.2 Experimental Setup
Model Series Dataset GSM8K Math GPQA Text-Avg MMStar MathVista MathVision MM-Avg Avg
@@ Appendix D Comprehensive Ablation Studies
MMStar MathVista MathVision MM-Avg GSM8K MATH GPQA Text-Avg

### v3

@@ 80. Guidelines:
Appendix A Implementation Details of MindGYM
A.1 Detailed Description of Meta-topics
  • Historical Events: Analyzing past events and their consequences.
  • Economic Principles: Studying resource allocation and valuation.
  • Psychological Insights: Exploring human behavior and cognition.
  • Temporal: Reasoning across sequential or causal timelines.
A.2 Utilized Prompts
A.2.1 Stream of Consciousness
A.2.2 Prompts for Background Generation
Write a background document of approximately 150-200 words. The document should describe a scenario or context that includes interconnected details, suitable for reasoning tasks. The document should focus on the reasoning category: {category}. The document should be rich enough to support multi-hop reasoning in Chinese.
A.2.3 Prompts for Seed Single-Hop Question Synthesis
Based on the background document provided, generate up to 5 logically connected sub-questions. The relationship between these sub-questions should belong to a single category: - Bridging: requiring connecting facts or pieces of information from the document. - Comparison: involving comparing two or more elements described in the document. - Temporal: requiring reasoning about the order or timing of events. Clearly state the relationship category that links all sub-questions in Chinese:
Based on the provided image, original question, and original answer, generate up to 5 logically connected sub-questions. The relationship between these sub-questions should belong to one of the following categories: - Visual-Textual Alignment: Requiring alignment between visual (e.g., images, charts) and textual information. - Spatial Reasoning: Involving spatial relationships or geometric layouts. - Causal Inference: Requiring reasoning about cause-and-effect relationships. - Contextual Synthesis: Requiring synthesis of information across multiple modalities (e.g., text, images, charts). Clearly state the relationship category that links all sub-questions in Chinese:
A.2.4 Prompts for Challenging Multi-Hop QA Synthesis
Combine the sub-questions into a single, complex multi-hop question. The question should require reasoning across the sub-questions and synthesizing information from the background document. Then, provide a detailed answer to the multi-hop question, ensuring it is consistent with the background document and sub-questions. Synthesize a multi-hop question and its answer based on the above sub-questions and background document in Chinese. Please start Qwen thinking and return the thinking process:
Combine the sub-questions into a single, complex multi-hop question. The question should require reasoning across the sub-questions and synthesizing information from the image and original question. Then, provide a detailed answer to the multi-hop question, ensuring it is consistent with the image and sub-questions. Synthesize a multi-hop question and its answer based on the above sub-questions, original image, and original question in Chinese. Please start Qwen thinking and return the thinking process:
A.2.5 Prompts for GPT-based Scorer
As a mathematical problem solver, please strictly compare and analyze the predictions of the two models: [Problem description] {question}
[Original model prediction] {pred_raw}
[Fine-tuned model prediction] {pred_finetuned}
Evaluation requirements: Give the scores of raw and finetuned respectively
1. Reasoning depth score (0-3 points): - 3 points: multi-step derivation with verification - 2 points: complete derivation steps - 1 point: simple calculation steps - 0 points: no derivation process
2. Reasoning breadth score (0-3): - 3 points: Explores multiple valid methods/angles, justifies optimal choice - 2 points: Mentions alternative approaches briefly but focuses on one - 1 points: Suggests another method without analysis - 0 points: Single approach with no alternatives considered
3. Comparative analysis: use bullet points to list the main improvements/degradations Please return a JSON that strictly follows this format: "accuracy_raw": 0-2, "accuracy_finetuned": 0-2, "depth_raw": 0-3, "depth_finetuned": 0-3, "comparison": ["Point 1", "Point 2"]
A.2.6 Prompts for Flux Data Sythesis
We use the currently best text-to-image model Flux to generate images from text, and compare it with data based on ScienceQA and OKVQA as image anchors. The process involves two main steps:
Step 1: Generating Structured Descriptions.
We first produce concise textual descriptions that serve as prompts for image generation. These descriptions are categorized into eight distinct educational domains to ensure coverage of diverse reasoning scenarios:
  • Geometry: basic geometric shapes with labeled lengths and angles.
  • Physics: simple mechanics such as pulleys, inclined planes, or force diagrams.
  • Chemistry: molecular structures or reaction schemes.
  • Math Word Problem: real-world scenes involving objects and quantities.
  • Logic Diagram: flowcharts or condition-based logical structures.
  • Statistics Chart: visualizations such as bar charts, pie charts, or line graphs.
  • Timeline or History Map: event timelines or migration maps.
  • Circuit Diagram: basic electronic circuits with labeled components.
Each category follows a unified generation prompt to ensure clarity and reliability: For each of the following categories, generate a simple and reliable prompt for a text-to-image generation model. The goal is to create educational images with accurate content, clear structure, and moderate visual complexity. The generated prompt should describe a plausible, textbook-style diagram with correct labels or values, suitable for multi-step reasoning (e.g. in math, science, or logic). The content should avoid errors, and not be too crowded or complex. Each prompt should be concise, standalone, and directly usable for image generation — do not include explanations or extra output. Use this category {category} and its description {description}:
Step 2: Image Synthesis via Flux.
The resulting category-specific prompts are then input into Flux, which generates corresponding images with consistent style and controlled complexity. Each synthesized image is paired with its associated textual reasoning data to form complete multimodal QA samples.
This two-step pipeline—structured prompt generation followed by Flux-based synthesis—ensures both semantic precision and visual clarity, enabling the dataset to support reasoning across textual and visual modalities.
Appendix B Detailed Experimental Setup
B.1 Benchmarks
@@
Given the extensive evaluation tasks, employing the complete evaluation sets would incur substantial time expenditure. To expedite the evaluation process while ensuring fairness and accuracy, we adopt the eval_limiteval\_limit parameter from ms-swift, configuring GPQA to 300 samples, while retaining the original sample sizes for the remaining benchmark datasets. All experiments were conducted using this consistent setup to ensure the fairness of the experiments.
B.2 Training and Evaluation Details
@@ C.2 Experimental Setup
Model Series Dataset GSM8K Math GPQA Text-Avg MMStar MathVista MathVision $M$-Avg Avg
@@ Appendix D Comprehensive Ablation Studies
MMStar MathVista MathVision $M$-Avg GSM8K MATH GPQA Text-Avg
@@
D.1 Complete Experimental Results
D.2 Comparison with Recent Proprietary Systems
Table 12: MathVision leaderboard comparison.
Model Score
GPT-4.5 47.3
Gemini-2 Flash 41.3
Qwen2.5-VL-72B-MindGym 39.47
Kimi k1.5 38.6
Qwen2.5-VL-72B (baseline) 38.10
Claude 3.5 Sonnet 37.99
Kimi-VL-A3B-Thinking 36.8
To situate MindGym’s data-driven improvements in context, we compare the fine-tuned Qwen2.5-VL-72B-MindGym model with leading proprietary reasoning systems on the MathVision leaderboard. While MindGym is not a new architecture but a synthesis framework, it delivers a consistent +1.37% gain over the original Qwen2.5-VL model, narrowing the gap to top commercial models. Closed-source systems such as o3 and o4 remain inaccessible for fine-tuning, but MindGym effectively enhances open-source models under comparable evaluation settings.
D.3 Comparison with MathFusion
MindGYM is conceptually related to MathFusion [29] in that both approaches leverage synthetic data to improve reasoning capabilities. However, MindGYM extends this idea by integrating multimodal and cognitively structured synthesis, covering a wider range of domains beyond purely mathematical tasks. In particular, the Thinking Claude framework embeds structured reasoning trajectories into both questions and answers, providing richer supervision for reasoning-aware fine-tuning.
To illustrate the difference, we compared a MathFusion-style synthesis with MindGYM on the same evaluation benchmarks:
Table 13: Comparison between MathFusion-style synthesis and MindGYM.
Method GSM8K Math GPQA text-avg MMStar MathVista MathVision $M$-avg Overall Avg
MathFusion-style 84.31 69.1 29.0 60.80 61.43 67.2 24.97 51.20 56.00
MindGYM 84.08 68.4 31.33 61.27 64.33 70.3 28.62 54.42 57.84
As shown in Table 13, MindGYM outperforms the MathFusion-style baseline on multimodal and general reasoning tasks, demonstrating the effectiveness of combining cognitive structure with multimodal synthesis.